OpenStudy (anonymous):

Multiply and express the product in simplest form. 15x^2-y times 4y^2/xy A.60x^2/y^2 B.60x C.15xy^3/4y^2 D.15x^3y/4y^3

OpenStudy (anonymous):

@davidmack

OpenStudy (anonymous):

help me please!

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\large (15x^2-y)\times \dfrac{ 4y^2}{xy}\)

ganeshie8 (ganeshie8):

like that ?

ganeshie8 (ganeshie8):

could u take snapshot and post if psble... ? :)

ganeshie8 (ganeshie8):

ty :)

ganeshie8 (ganeshie8):

\(\large \dfrac{ 15x^2}{y} \times \dfrac{ 4y^2}{xy}\)

ganeshie8 (ganeshie8):

first multiply the constants : 15*4 wat do u get ?

OpenStudy (anonymous):

60

ganeshie8 (ganeshie8):

\(\large \dfrac{ 15x^2}{y} \times \dfrac{ 4y^2}{xy}\) \(\large \dfrac{ 60x^2}{y} \times \dfrac{ y^2}{xy}\)

ganeshie8 (ganeshie8):

yes :)

ganeshie8 (ganeshie8):

bottom you have two y's, and top you have y^2. can we cancel them both ?

OpenStudy (anonymous):

okay..umm...yes?

ganeshie8 (ganeshie8):

\(\large \dfrac{ 15x^2}{y} \times \dfrac{ 4y^2}{xy}\) \(\large \dfrac{ 60x^2}{y} \times \dfrac{ y^2}{xy}\) \(\large 60x^2 \times \dfrac{1 }{x}\)

ganeshie8 (ganeshie8):

lastly cancel one x also

ganeshie8 (ganeshie8):

\(\large 60x^2 \times \dfrac{1 }{x}\) \(\large 60x\)

ganeshie8 (ganeshie8):

see if that makes more or less sense.. .

OpenStudy (anonymous):

omg thank you soo much and yes it does! :)

ganeshie8 (ganeshie8):

good to hear :) u wlc !